# thanks: Corresponding author

### v1

@@ Abstract
We present a perturbative triples correction to the relativistic quadratic unitary coupled cluster singles and doubles (qUCCSD) method, denoted as qUCCSD[T]. The method builds upon the Hermitian structure of the unitary ansatz and employs a many-body perturbation theory framework to consistently include the effects of triple excitations without the need for computationally intensive iterative procedures. Relativistic effects are incorporated using the exact two-component atomic mean-field (X2CAMF) Hamiltonian, and the computational cost is further reduced through the frozen natural spinor (FNS) and Cholesky decomposition (CD) approximations. Benchmark results demonstrate that qUCCSD[T] outperforms previously proposed triples corrections to the unitary coupled cluster method in the clasical computing regime and yields excellent agreement with experimental data and Full CI benchmarks. Specifically, the method shows high accuracy in computing bond dissociation enthalpies, molecular geometries, vibrational frequencies, ionization potentials, and electron affinities of heavy-element-containing systems.
@@ I Introduction
The accurate treatment of electron correlation remains one of the central challenges in quantum chemistry and chemical physics. The coupled cluster (CC) method[1, 2], which uses an exponential parametrization of the wave function, has emerged as one of the most accurate and systematically improvable methods. The exponential parametrization guarantees the size extensivity of the energy, even in finite truncation of the cluster operator. The coupled cluster method is generally used in singles and doubles approximation of the cluster operator (CCSD). One can systematically improve the CCSD method by including operators of higher excitation ranks[3, 4, 5, 6]. The CCSD(T) method[7, 8] is particularly popular due to its inherent balance of computational cost and accuracy, and is considered the ”gold standard” of quantum chemistry.
@@
In addition to the standard coupled cluster method, the unitary coupled cluster (UCC) variants[9, 10, 11, 12, 13] have attracted considerable attention[14] owing to their Hermitian nature and potential advantages for both classical and quantum computing platforms. The unitary coupled cluster method tends to converge[15] to Full CI (FCI) results when higher excitation operators are considered. However, unlike the standard coupled cluster method, there is no natural truncation in Hamiltonian cluster commutator expansion in the UCC method, and it scales as FCI even with a finite truncation of the cluster operator. Consequently, one cannot solve the unitary coupled cluster method on classical computers without evoking additional approximations. One of the earliest adopted approximations relies on the perturbation theory-based truncation[9] of the UCC energy functionals. Perturbational approximation-based unitary coupled cluster method (UCC(n), where n denotes the perturbation order) has been extended to triples and quadrupole excitation[16] operators, too. Bartlett and co-workers[17] have recently reported a non-iterative triples correction scheme to the standard and perturbational approximation-based unitary coupled cluster method. However, the approximate unitary CC method based on the lower order of perturbation expansion often gives inferior results[18, 19] compared to the standard coupled cluster method.
@@
An alternative approach for the truncation of the unitary coupled cluster method is based on the commutator rank[13, 20]. Among the various commutator-based truncation schemes of the unitary coupled cluster method, the quadratic unitary coupled cluster method(qUCC)[21, 22] has emerged as one giving the best compromise[23] between computational cost and accuracy. The quadratic unitary coupled cluster method is generally used in the singles and doubles approximation(qUCCSD)[21] and has been implemented for ground and excited state energies[21, 22, 24]. The natural orbital based low-cost implementation of qUCCSD and extension to the relativistic domain has also been achieved[24]. Yet, for systems where dynamic correlation is significant, particularly heavy-element systems, contributions from triple excitations are crucial for achieving chemical accuracy in addition to the relativistic effect[25]. Fully iterative inclusion of triple excitations within the quadratic unitary coupled cluster is computationally prohibitive for all but the smallest systems, motivating the need for efficient perturbative treatments. At the same time, cost-effective treatment of the relativistic effect is necessary to make the method practical for real-life systems. In this work, we develop a perturbative triple correction to the relativistic quadratic Unitary Coupled Cluster Singles and Doubles (qUCCSD) method, denoted as qUCCSD[T]. The exact two-component Hamiltonian with the atomic mean field treatment of the spin-orbit coupling effect (X2AMF)[26, 27, 28] has been used to introduce the relativistic effect. Building upon the underlying structure of qUCCSD, we derive a consistent and computationally tractable perturbative framework to account for triple excitations. The paper is organized as follows: Section II presents the theoretical development of the perturbative triple correction within the relativistic quadratic unitary coupled cluster framework. Section III provides benchmarking results on a set of molecular systems, followed by conclusions in Section IV.
@@ II.1 Relativistic quadratic unitary coupled cluster theory
where $i$, $j$, $k$, $l$ and $a$, $b$, $c$, $d$ symbols represent the occupied and virtual spinors, respectively. Inclusion of
@@
gives rise to the UCCSDT method. The cluster amplitudes($\sigma$) are determined by simultaneously solving a system of nonlinear equations
@@
A notable drawback of employing an anti-hermitian cluster operator in the unitary coupled cluster framework is that the Baker-Campbell-Hausdorff (BCH) expansion
@@
does not naturally truncate at finite order. Consequently, the practical application of UCC requires an artificial truncation of this infinite series. However, a significant challenge lies in the lack of a universally preferred or unique scheme for this truncation. There are multiple strategies for truncating the BCH expansion in Eq. 11. A popular approach involves applying arguments from many-body perturbation theory (MBPT)[18, 19]. Alternatively, one can truncate the BCH expansion based on the depth of the commutator rank. One of the attractive schemes for commutator-based truncation of the similarity transformed Hamiltonian in the unitary coupled cluster method[20] is based on the Bernoulli expansion[29, 30]. In this approach, the $\hat{H}$ is partitioned into the Fock operator ($\hat{F}$) and a fluctuation potential ($\hat{V}$) as
@@
The Fock operator is block diagonal and rank-conserving operator for the canonical Hartree-Fock method
@@
can be further separated into the non-diagonal ($\hat{V}_{N}$) part which consist of pure excitation and de-excitation operator and the ”rest” part ($\hat{V}_{R}$). Now, $\bar{H}$ can be expanded using Bernoulli numbers
$e^{-\sigma}\hat{H}e^{\sigma}=\bar{H}_{0}+\bar{H}_{1}+\bar{H_{2}}+\bar{H_{3}}\dots$ (15)
@@
$\bar{H}_{0}$ $=$ $\hat{F}+\hat{V}$ (16)
$\bar{H}_{1}$ $=$ $[\hat{F},\hat{\sigma}]+\frac{1}{2}[\hat{V},\hat{\sigma}]+\frac{1}{2}[\hat{V}_{R},\hat{\sigma}$] (17)
$\bar{H}_{2}$ $=$ $\frac{1}{12}[[\hat{V}_{N},\hat{\sigma}],\hat{\sigma}]+\frac{1}{4}[[\hat{V},\hat{\sigma}]_{R},\hat{\sigma}$] (18)
@@
$\bar{H}_{3}$ $=$ $\frac{1}{24}[[[\hat{V}_{N},\hat{\sigma}],\hat{\sigma}]_{R},\hat{\sigma}]+\frac{1}{8}[[[[\hat{V}_{R},\hat{\sigma}]_{R},\hat{\sigma}]_{R},\hat{\sigma}$] (19)
@@
Such a structure not only simplifies the formulation but also provides a rigorous and efficient foundation for developing non-perturbative approximations within the unitary coupled cluster framework. One can derive an approximation to UCC method by taking the commutator up to a particular rank. For example, taking terms up to a $\bar{H}_{3}$ in the energy and $\bar{H}_{2}$ in the amplitude equation leads to a quadratic unitary coupled cluster approximation[21].
$E_{Gr}^{\text{qUCC}}=E^{\text{HF}}+\langle\phi_{0}|\bar{H}^{1}|\phi_{0}\rangle+\langle\phi_{0}|\bar{H}^{2}|\phi_{0}\rangle$
$+\langle\phi_{0}|\bar{H}^{3}|\phi_{0}\rangle$ (20)
$\langle\phi_{i}^{a}|\bar{H}_{1}+\bar{H}_{2}|\phi_{0}\rangle=0$
$\langle\phi_{ij}^{ab}|\bar{H}_{0}+\bar{H}_{1}+\bar{H}_{2}|\phi_{0}\rangle=0$
$\langle\phi_{ijk}^{abc}|\bar{H}_{1}+\bar{H}_{2}|\phi_{0}\rangle=0$
@@ II.2 Perturbative triples correction to qUCCSD method
One of the most advantageous features of the qUCC method is that one can directly derive the perturbative triples correction from the perturbation-based truncation of the qUCCSDT energy expression in equation 20. It is conceptually more straightforward than the CCSD(T) method in the standard coupled cluster method, where a rigorous derivation of the perturbative triples correction requires switching back and forth between projection and variational energy functional[31, 16, 32]. One can formulate the perturbative triples correction scheme to the qUCC method using many-body perturbation theory(MBPT), assuming a canonical Hartree-Fock (HF) reference state. Within this framework, the Fock operator $\hat{F}$ and the fluctuation operator $\hat{V}$ contribute at zeroth and first order in MBPT, respectively. The excitation operator $\hat{\sigma}_{n}$ typically emerges at the (n-1) order, with the exception of $\hat{\sigma}_{1}$, which contributes the second order. The qUCCSD method is complete up to fourth order in energy within singles and doubles truncation of the cluster operator[21]. Additional terms arise due to the three-body operator, which will make the energy expression complete up to the fourth order are
@@
$\quad\phantom{-}\frac{1}{8}\langle kj||ai\rangle\sigma(2)_{abc}^{lkj}\sigma_{cb}^{il}+\frac{1}{8}\langle ia||bc\rangle\sigma_{da}^{jk}\sigma(2)_{bcd}^{kji}$
$\quad-\frac{1}{8}\langle ka||ij\rangle\sigma(2)_{cba}^{jil}\sigma_{bc}^{lk}-\frac{1}{8}\langle ba||ci\rangle\sigma(2)_{dab}^{ijk}\sigma_{cd}^{kj}$ (22)
Where $\sigma_{a}^{i}$ and $\sigma_{ab}^{ij}$ are converged qUCCSD singles and doubles amplitudes. The $\sigma(2)^{ijk}_{abc}$ is the second-order triples amplitude
$\sigma(2)_{abc}^{ijk}=\frac{s_{abc}^{ijk}}{\epsilon_{a}+\epsilon_{b}+\epsilon_{c}-\epsilon_{i}-\epsilon_{j}-\epsilon_{k}}$ (23)
@@ II.3 Exact Two-component Hamiltonian with the atomic mean field approximation
One of the most rigorous ways to include the relativistic effect is to use a four-component Dirac Coulomb(4c-DC) Hamiltonian[33]. The high computational cost associated with 4c-DC Hamiltonian restricts its applicability beyond atoms and small molecules. One of the practical ways to reduce the computational cost of relativistic calculations is to employ two-component Hamiltonians. Among the various flavors of two-component Hamiltonians available[34, 35, 36, 37, 38, 39, 40, 41] , we are going to use the exact two-component atomic mean field(X2CAMF) approach[26, 27, 28]. The 4c-DC Hamiltonian can be defined as
@@
Within the no-pair approximation[42, 33], the summation in the above equation is confined only to the positive-energy spinors. The indices $p$,$q$, $r$, and $s$ represented the positive-energy four-component spinors and the creation and the second quantized creation and annihilation operators are denoted by $\hat{a}_{p}^{\dagger}$, $\hat{a}_{q}^{\dagger}$ and $\hat{a}_{r}$, $\hat{a}_{s}$ respectively. In the spin separation scheme, the matrix elements of two-electron interaction can be split into their spin-free (SF) and spin-dependent (SD) parts
@@
Taking advantage of the localized nature of the spin-orbit interaction, the spin-dependent term can be approximated by the atomic mean field(AMF) approximation and the spin-free term can be approximated by the non-relativistic two-electron integrals
@@
One of the most prominent advantages of this approach is that the required two electrons are only non-relativistic. Therefore, one can use the already well-established techniques for the efficient treatment of non-relativistic two-electron integrals in the relativistic calculations. In the present work, the two-electron integrals are treated using the Cholesky decomposition(CD) technique[43, 44, 45, 46].

### v2

@@ Abstract
We present a non-iterative triples correction to the relativistic quadratic unitary coupled cluster singles and doubles (qUCCSD) method, denoted as qUCCSD[T]. The method builds upon the Hermitian structure of the similarity transformed Hamiltonian in unitary coupled cluster method and can be derived using perturbational truncation to unitary coupled cluster energy functional. Relativistic effects are incorporated using the exact two-component atomic mean-field (X2CAMF) Hamiltonian, and the computational cost is further reduced through the frozen natural spinor (FNS) and Cholesky decomposition (CD) approximations. Benchmark results demonstrate that qUCCSD[T] outperforms previously proposed triples corrections to the unitary coupled cluster method and yields excellent agreement with experimental data and Full CI results. Additionally, the method shows high accuracy in computing the bond dissociation enthalpies, molecular geometries, vibrational frequencies, ionization potentials, and electron affinities of heavy-element-containing systems. The new qUCCSD[T] method is competitive to popular CCSD(T) even on a classical computer.
@@ I Introduction
The accurate treatment of electron correlation remains one of the central challenges in quantum chemistry and chemical physics. The coupled cluster (CC) method,[1, 2] which uses an exponential parametrization of the wave function, has emerged as one of the most accurate and systematically improvable methods. The exponential parametrization guarantees the size extensivity of the energy, even in finite truncation of the cluster operator. The coupled cluster method is generally used in singles and doubles approximation of the cluster operator (CCSD). One can systematically improve the CCSD method by including operators of higher excitation ranks.[3, 4, 5, 6] The CCSD(T) method[7, 8] is particularly popular due to its inherent balance of computational cost and accuracy, and is considered the “gold standard” of quantum chemistry.
@@
In addition to the standard coupled cluster method, the unitary coupled cluster (UCC) variants[9, 10, 11, 12, 13] have attracted considerable attention[14] owing to their Hermitian nature and potential advantages for both classical and quantum computing platforms. The UCC method tends to converge[15] to Full CI (FCI) results when higher excitation operators are considered. However, unlike the standard CC method, there is no natural truncation in Hamiltonian cluster commutator expansion in the UCC method, and it scales as FCI even with a finite truncation of the cluster operator. Consequently, one cannot solve the UCC method on classical computers without evoking additional approximations. One of the earliest adopted approximations relies on the perturbation theory-based truncation[9] of the UCC energy functionals. Perturbational approximation-based unitary coupled cluster method (UCC(n)), where n denotes the perturbation order) has been extended to triples and quadrupole excitation[16] operators, too. Bartlett and co-workers[17] have recently reported a non-iterative triples correction scheme to the standard and perturbational approximation-based UCC method. However, the approximate UCC method based on the lower order of perturbation expansion often gives inferior results[18, 19] compared to the standard CC method.
@@
An alternative approach for the truncation of the UCC method is based on the commutator rank.[13, 20] Among the various commutator-based truncation schemes of the UCC method, the quadratic unitary coupled cluster method(qUCC)[21, 22] has emerged as one giving the best compromise[23] between computational cost and accuracy. The qUCC method is generally used in the singles and doubles approximation(qUCCSD)[21] and has been implemented for ground and excited state energies.[21, 22, 24] The natural orbital based low-cost implementation of qUCCSD and extension to the relativistic domain has also been achieved.[24] Yet, for systems where dynamic correlation is significant, particularly heavy-element systems, contributions from triple excitations are crucial for achieving chemical accuracy in addition to the relativistic effect[25]. Fully iterative inclusion of triple excitations within the qUCC is computationally prohibitive for all but the smallest systems, motivating the need for efficient perturbative treatments. At the same time, cost-effective treatment of the relativistic effect is necessary to make the method practical for real-life systems. In this work, we develop a perturbative triple correction to the relativistic qUCCSD method, denoted as qUCCSD[T]. The exact two-component Hamiltonian with the atomic mean field treatment of the spin-orbit coupling effect (X2CAMF)[26, 27, 28] has been used to introduce the relativistic effect. Building upon the underlying structure of qUCCSD, we derive a consistent and computationally tractable perturbative framework to account for triple excitations. The paper is organized as follows: Section II presents the theoretical development of the perturbative triple correction within the relativistic qUCC framework. Section III provides benchmarking results on a set of molecular systems, followed by conclusions in Section IV.
@@ II.1 Relativistic quadratic unitary coupled cluster theory
Inclusion of
@@
gives rise to the UCCSDT method. where $i$, $j$, $k$, $l$ denote the occupied spinors, while $a$, $b$, $c$, $d$ represent virtual spinors. The quantities $\sigma_{i}^{a}$, $\sigma_{ij}^{ab}$ and $\sigma_{ijk}^{abc}$ are the singles, doubles and triples amplitudes, respectively. These cluster amplitudes($\sigma$) are determined by simultaneously solving a system of nonlinear equations
@@
A notable drawback of employing an anti-hermitian cluster operator in the UCC framework is that the Baker-Campbell-Hausdorff (BCH) expansion
@@
does not naturally truncate at finite order. Consequently, the practical application of UCC requires an artificial truncation of this infinite series. However, a significant challenge lies in the lack of a universally preferred or unique scheme for this truncation. There are multiple strategies for truncating the BCH expansion in Eq. 11. A popular approach involves applying arguments from many-body perturbation theory (MBPT).[18, 19] Alternatively, one can truncate the BCH expansion based on the depth of the commutator rank. One of the attractive schemes for commutator-based truncation of the similarity transformed Hamiltonian in the UCC method[20] is based on the Bernoulli expansion.[29, 30] In this approach, the $\hat{H}$ is partitioned into the Fock operator ($\hat{F}$) and a fluctuation potential ($\hat{V}$) as
@@
The Fock operator is block diagonal and rank-conserving operator for the canonical Dirac Hartree-Fock method
@@
can be further separated into the non-diagonal ($\hat{V}_{N}$) part which consist of pure excitation and de-excitation operator and the “rest” part ($\hat{V}_{R}$). Now, $\bar{H}$ can be expanded using Bernoulli numbers
$\bar{H}=\bar{H}^{0}+\bar{H}^{1}+\bar{H^{2}}+\bar{H^{3}}\dots$ (15)
@@
$\bar{H}^{0}$ $=$ $\hat{F}+\hat{V}$ (16)
$\bar{H}^{1}$ $=$ $[\hat{F},\hat{\sigma}]+\frac{1}{2}[\hat{V},\hat{\sigma}]+\frac{1}{2}[\hat{V}_{R},\hat{\sigma}$] (17)
$\bar{H}^{2}$ $=$ $\frac{1}{12}[[\hat{V}_{N},\hat{\sigma}],\hat{\sigma}]+\frac{1}{4}[[\hat{V},\hat{\sigma}]_{R},\hat{\sigma}$] (18)
@@
$\bar{H}^{3}$ $=$ $\frac{1}{24}[[[\hat{V}_{N},\hat{\sigma}],\hat{\sigma}]_{R},\hat{\sigma}]+\frac{1}{8}[[[[\hat{V}_{R},\hat{\sigma}]_{R},\hat{\sigma}]_{R},\hat{\sigma}$] (19)
@@
Such a structure not only simplifies the formulation but also provides a rigorous and efficient foundation for developing non-perturbative approximations within the UCC framework. One can derive an approximation to UCC method by taking the commutator up to a particular rank. For example, taking terms up to a $\bar{H}^{3}$ in the energy and $\bar{H}^{2}$ in the amplitude equation leads to a quadratic unitary coupled cluster approximation.[21]
$E_{Gr}^{\text{qUCC}}=\langle\Phi_{0}|\bar{H}^{0}+\bar{H}^{1}+\bar{H}^{2}+\bar{H}^{3}|\Phi_{0}\rangle$ (20)
$\langle\Phi_{i}^{a}|\bar{H}^{0}+\bar{H}^{1}+\bar{H}^{2}|\Phi_{0}\rangle=0$
$\langle\Phi_{ij}^{ab}|\bar{H}^{0}+\bar{H}^{1}+\bar{H}^{2}|\Phi_{0}\rangle=0$
$\langle\Phi_{ijk}^{abc}|\bar{H}^{0}+\bar{H}^{1}+\bar{H}^{2}|\Phi_{0}\rangle=0$
@@ II.2 Perturbative triples correction to qUCCSD method
One of the most advantageous features of the qUCC method is that one can directly derive the perturbative triples correction from the perturbation-based truncation of the qUCCSDT energy expression in equation 20. It is conceptually more straightforward than the CCSD(T) method in the standard coupled cluster method, where a rigorous derivation of the perturbative triples correction requires switching back and forth between projection and variational energy functional.[8, 16, 31] One can formulate the perturbative triples correction scheme to the qUCC method using MBPT, assuming a canonical Hartree-Fock (HF) reference state. Within this framework, the Fock operator $\hat{F}$ and the fluctuation operator $\hat{V}$ contribute at zeroth and first order in MBPT, respectively. The excitation operator $\hat{\sigma}_{n}$ typically emerges at the (n-1) order, with the exception of $\hat{\sigma}_{1}$, which contributes the second order. The qUCCSD method is complete up to fourth order in energy within singles and doubles truncation of the cluster operator.[21] Additional terms arise due to the three-body operator, which will make the energy expression complete up to the fourth order are
@@
$\quad\phantom{-}\frac{1}{8}\langle kj||ai\rangle\sigma_{abc}^{lkj(2)}\sigma_{cb}^{il}+\frac{1}{8}\langle ia||bc\rangle\sigma_{da}^{jk}\sigma_{bcd}^{kji(2)}$
$\quad-\frac{1}{8}\langle ka||ij\rangle\sigma_{cba}^{jil(2)}\sigma_{bc}^{lk}-\frac{1}{8}\langle ba||ci\rangle\sigma_{dab}^{ijk(2)}\sigma_{cd}^{kj}$ (22)
Where $\sigma_{a}^{i}$ and $\sigma_{ab}^{ij}$ are converged qUCCSD singles and doubles amplitudes. The $\sigma^{ijk(2)}_{abc}$ is the second-order triples amplitude
$\sigma_{abc}^{ijk(2)}=\frac{s_{abc}^{ijk}}{\epsilon_{a}+\epsilon_{b}+\epsilon_{c}-\epsilon_{i}-\epsilon_{j}-\epsilon_{k}}$ (23)
@@ II.3 Exact Two-component Hamiltonian with the atomic mean field approximation
One of the most rigorous ways to include the relativistic effect is to use a four-component Dirac Coulomb(4c-DC) Hamiltonian.[32] The high computational cost associated with 4c-DC Hamiltonian restricts its applicability beyond small to medium-sized systems. One of the practical ways to reduce the computational cost of relativistic calculations is to employ two-component Hamiltonians. Among the various flavors of two-component Hamiltonians available,[33, 34, 35, 36, 37, 38, 39, 32] we are going to use the X2CAMF approach.[26, 27, 28] The 4c-DC Hamiltonian can be defined as
@@
Within the no-pair approximation,[40, 32] the summation in the above equation is confined only to the positive-energy spinors. The indices $p$,$q$, $r$, and $s$ represented the positive-energy four-component spinors and the second quantized creation and annihilation operators are denoted by $\hat{a}_{p}^{\dagger}$, $\hat{a}_{q}^{\dagger}$ and $\hat{a}_{r}$, $\hat{a}_{s}$ respectively. In the spin separation scheme, the matrix elements of two-electron interaction can be split into their spin-free (SF) and spin-dependent (SD) parts
@@
Taking advantage of the localized nature of the spin-orbit interaction, the spin-dependent term can be approximated by the atomic mean field(AMF) approximation and the spin-free term can be approximated by the non-relativistic (NR) two-electron integrals
@@
One of the most prominent advantages of this approach is that the required two electrons are only non-relativistic. Therefore, one can use the already well-established techniques for the efficient treatment of non-relativistic two-electron integrals in the relativistic calculations. In the present work, the two-electron integrals are treated using the Cholesky decomposition(CD) technique.[41, 42, 43, 44]